### v2

@@ 4. Other graphs characterizing dendricity
Let $X$ be the Thue-Morse shift space over $\{0,1\}$. The first few graphs $G^{L}_{n}(X)$ are given in Figure 2.
@@
Let $X$ be a shift space and $v\in\mathcal{L}_{n}(X)$. The graph $\mathcal{E}_{X}(v)$ contains the path $(a_{1}^{L},b_{1}^{R},a_{2}^{L},\dots,b_{k}^{R},a_{k+1}^{L})$ if and only if $G^{L}_{n+1}(X)$ contains the path
$a_{1}\xrightarrow{vb_{1}}a_{2}\dots\xrightarrow{vb_{k}}a_{k+1}.$
Symmetrically, the graph $\mathcal{E}_{X}(v)$ contains the path $(a_{1}^{R},b_{1}^{L},a_{2}^{R},\dots,b_{k}^{L},a_{k+1}^{R})$ if and only if $G^{R}_{n+1}(X)$ contains the path
$a_{1}\xrightarrow{b_{1}v}a_{2}\dots\xrightarrow{b_{k}v}a_{k+1}.$
@@
Let $X$ be a shift, $N\geq 0$ and $C\subset\mathcal{A}$. If the graph $\mathcal{E}_{X}(v)$ is acyclic for all $v\in\mathcal{L}_{<N}(X)$, then the following properties are equivalent.
@@
Let us show the first set of equivalences. We denote by $H_{n}$ the subgraph of $G^{L}_{n}(X)$ generated by the vertices in $C$. Assume that the first property is satisfied. We prove that $H_{n}$ is connected by induction on $n$. The graph $G^{L}_{0}(X)$ is a complete graph thus any subgraph is connected. If $H_{n}$ is connected, then, to prove that $H_{n+1}$ is also connected, it suffices to show that any two vertices $a,b\in C$ that were connected by an edge in $H_{n}$ are connected by a path in $H_{n+1}$. Let $a,b\in C$ be two such vertices and $v$ be the label of an edge between them. As $a$ and $b$ are left extensions of $v$, $a^{L}$ and $b^{L}$ are connected in $\mathcal{E}_{X}(v)$ by a path avoiding vertices $c^{L}$, $c\notin C$. The vertices $a$ and $b$ are thus connected by a path in $H_{n+1}$ by Lemma 24. Notice that the acyclic hypothesis is not needed for this implication.
@@ 5. Stabilization and eventually dendric shifts
To make the distinction clearer, we will talk about multi-graphs with colored edges instead of labeled edges. We can naturally extend the notion of acyclicity for the labeling to acyclicity for the coloring.
@@
A multi-graph $G$ with colored edges and with set of vertices $V$ is multi-clique if there exist subsets $C_{1},\dots,C_{k}$ of $V$ such that the set of edges is the union of the sets of colored edges $\left((C_{i}\times C_{i})\setminus\operatorname{diag}(C_{i})\right)\times\{c_{i}\}$, $i\leq k$, where $c_{i}$ is the color of the edges and $c_{1},\dots,c_{k}$ are distinct colors. The multi-clique $G$ is then denoted $G(\{C_{1},\dots,C_{k}\})$.
@@
If $C=\{0,1,2,3\}$, $C_{1}=\{0,1\}$, $C_{2}=\{1,2,3\}$, $C_{3}=\{0,3\}$, the multi-clique $G(\{C_{1},C_{2},C_{3}\})$ is represented in Figure 4. It is not acyclic from the coloring because of the cycle $0,1,3,0$.
@@
Another way to see the graph $G^{L}(X)$ is to consider the right infinite words and their left extensions. More precisely, for a right infinite word $x$, we denote by $E^{L}_{X}(x)$ the set of letters $a$ such that $\mathcal{L}(ax)\subset\mathcal{L}(X)$. We then have
@@
Let $\sigma$ be the morphism defined by $\sigma(0)=0110$ and $\sigma(1)=011$. It is a return morphism for $01$. Let $X$ be the image under $\sigma$ of a Sturmian shift. It is not a dendric shift space since the graph $\mathcal{E}_{X}(\varepsilon)$ contains a cycle but, by Remark 15, $X$ is eventually dendric with threshold at most 4. We can then show that the only words which are not dendric are $\varepsilon$ and $1$ thus $G^{L}(X)=G^{L}_{2}(X)$ and $G^{R}(X)=G^{R}_{2}(X)$. These graphs are acyclic for the coloring and connected since $01$ (resp., $10$) is the only left (resp., right) special word of length 2.
@@ 6. Action of morphisms on $G^{L}(X)$ and $G^{R}(X)$
Let us prove the link between $G^{L}(X)$ and $G^{L}(Y)$. By Proposition 19, $Y$ is also eventually dendric. If we write $m=\max\{|\sigma(a)|:a\in\mathcal{A}\}$, let $N\in\mathbb{N}$ (resp., $M\in\mathbb{N}$) be large enough so that $N$ (resp., $M-m$) satisfies the condition 2 of Proposition 17 for the shift $X$ (resp., $Y$). Let us also assume that the antecedent of any word $u\in\mathcal{L}(Y)$ of length at least $M$ is of length at least $N$. By Proposition 32, $G^{L}(X)$ (resp., $G^{L}(Y)$) is constructed with the left special factors of length $N$ (resp., $M$).
We define the application $f$ mapping any left special $u\in\mathcal{L}_{M}(Y)$ to the pair $(s_{u},v)\in\mathcal{T}^{L}(\sigma)\times\mathcal{L}_{N}(X)$ where $(s_{u},v_{u},p_{u})$ is the triplet associated with $u$ by Proposition 8 and $v$ is the prefix of length $N$ of $v_{u}$.
@@
$E^{L}_{Y}(u)\subset E^{L}_{Y}(s\sigma(v)w)=\varphi^{L}_{\sigma,s}(E^{L}_{X}(v))$
@@
The application $f$ is injective. Indeed, if $f(u)=f(u^{\prime})=(s,v)$, then $s_{u}=s=s_{u^{\prime}}$ and $v$ is a prefix of both $v_{u}$ and $v_{u^{\prime}}$. However, as $u$ and $u^{\prime}$ are left special, $v_{u}$ and $v_{u^{\prime}}$ must be left special. As they have a common prefix of length $N$ and $N$ is chosen large enough, it means that one is prefix of the other. Let us assume that $v_{u}$ is a prefix of $v_{u^{\prime}}$. Then $s\sigma(v_{u})w$ is a prefix of both $u$ and $u^{\prime}$. Since $u$ is of length $M$, $s\sigma(v_{u})w$ is a left special factor of length at least $M-m$ and, as $M$ is large enough, it is prefix of a unique left special factor of length $M$. This proves that $u=u^{\prime}$ and that $f$ is injective.
@@
$E^{L}_{Y}(u)\supset E^{L}_{Y}(s\sigma(v^{\prime})w)=\varphi_{\sigma,s}^{L}(E^{L}_{X}(v^{\prime}))=\varphi_{\sigma,s}^{L}(E^{L}_{X}(v)).$
@@ 7.1. $S$-adic representations
As $\max_{a\in\mathcal{A}_{N}}|\sigma_{[n,N)}(a)|$ goes to infinity when $N$ increases, $\mathcal{L}^{(n)}({{\boldsymbol{\sigma}}})$ defines a non-empty shift space $X_{{\boldsymbol{\sigma}}}^{(n)}$. More precisely, $X_{{\boldsymbol{\sigma}}}^{(n)}$ is the set of points $x\in\mathcal{A}_{n}^{\mathbb{Z}}$ such that $\mathcal{L}(x)\subset\mathcal{L}^{(n)}({{\boldsymbol{\sigma}}})$. Note that it may happen that $\mathcal{L}(X_{{\boldsymbol{\sigma}}}^{(n)})$ is strictly contained in $\mathcal{L}^{(n)}({{\boldsymbol{\sigma}}})$. Also observe that for all $n$, $X_{{\boldsymbol{\sigma}}}^{(n)}$ is the image of $X_{{\boldsymbol{\sigma}}}^{(n+1)}$ under $\sigma_{n}$.
We set $X_{{\boldsymbol{\sigma}}}=X_{{\boldsymbol{\sigma}}}^{(0)}$ and call ${\boldsymbol{\sigma}}$ an $S$-adic representation of $X_{{\boldsymbol{\sigma}}}$.
@@
We let $\mathcal{R}_{X}(w)$ denote the set of return words to $w$ in $X$ and we omit the subscript $X$ whenever it is clear from the context. The shift space $X$ being minimal, $\mathcal{R}(w)$ is always finite thus we write $R_{X}(w)=\{1,\dots,\operatorname{Card}(\mathcal{R}_{X}(w))\}$. A morphism $\sigma:{R(w)}^{*}\to\mathcal{A}^{*}$ is a coding morphism associated with $w$ if $\sigma(R(w))=\mathcal{R}(w)$. It is trivially a return morphism for $w$.
Let us consider the set $\mathcal{D}_{w}(X)=\{x\in{R(w)}^{\mathbb{Z}}\mid\sigma(x)\in X\}$. It is a minimal shift space, called the derived shift of $X$ (with respect to $w$). We now show that derivation of minimal shift spaces allows to build primitive $\mathcal{S}$-adic representations with return morphisms. We inductively define the sequences $(w_{n})_{n\geq 0}$, $(R_{n})_{n\geq 0}$, $(X_{n})_{n\geq 0}$ and $(\sigma_{n})_{n\geq 0}$ by
@@
Let $X$ be a minimal dendric shift over the alphabet $\mathcal{A}=\{1,\dots,d\}$. For any $w\in\mathcal{L}(X)$, $\mathcal{D}_{w}(X)$ is a minimal dendric shift over $\mathcal{A}$ and the coding morphism associated with $w$ is a tame automorphism of $F_{\mathcal{A}}$. As a consequence, if ${\boldsymbol{\sigma}}=(\sigma_{n})_{n\geq 0}$ is the primitive directive sequence of Theorem 42, then all morphisms $\sigma_{n}$ are tame dendric return morphisms.
@@
In particular, if $w_{n}$ is a letter, the morphism $\sigma_{n}$ is strongly left proper, i.e. the image of each letter begins with the letter $w_{n}$ which does not appear elsewhere. Moreover, we can assume that $w_{n}$ is a bispecial letter, as shown in the following results. This implies that $w_{n}\in\mathcal{T}^{R}(\sigma_{n})$ and $\varepsilon\in\mathcal{T}^{L}(\sigma_{n})$.
@@
First, notice that, because $\mathcal{E}_{X}(\varepsilon)$ is connected, there is always a left and a right special letter. Let $a_{1}$ be a right special letter. If it is bispecial, we can conclude. Otherwise, let $a_{2}$ be its unique left extension. Using Lemma 44, we know that $a_{2}\neq a_{1}$. In the graph $\mathcal{E}_{X}(\varepsilon)$, $a_{2}^{L}$ is then the only neighbour of the vertex $a_{1}^{R}$ but this graph is connected thus the vertex $a_{2}^{L}$ is of degree at least two and $a_{2}$ is a right special letter. Once again, if it is left special, we have found a bispecial letter and otherwise it has a unique left extension $a_{3}$. By Lemma 44, we must have $a_{3}\neq a_{1}$ and $a_{3}\neq a_{2}$. We can iterate the process to define $a_{4},a_{5},\dots$. However, since the alphabet is finite, it stops at some point, meaning that we have found a bispecial letter. ∎
@@ 7.2. Characterization via two graphs
We can deduce that a shift space is minimal dendric if and only if it has a primitive $S$-adic representation labeling an infinite path in the graphs built with the left (resp., right) valid triplets for dendric return morphisms.
We do not go into more details as we will prove that, instead of considering all the valid triplets, we can restrict ourselves to the case where the graphs are colored trees, i.e. all the edges have a different color and the corresponding uncolored graph is a tree. Note that a colored tree is trivially acyclic for the coloring and connected. We will often say that the colored graph itself is a tree and assimilate the colored and the uncolored version.
@@
  • Otherwise, $as$ and $s^{\prime}$ are not suffix comparable thus $\mathcal{A}^{L}_{\sigma,s^{\prime}}\cap C_{i}=\mathcal{A}^{L}_{\sigma,s^{\prime}}\cap E$, $\mathcal{A}^{L}_{\sigma,s^{\prime}}\cap D\subset\{c\}$ and no edge was lost from $G^{L}_{\sigma,s^{\prime}}$ to $H^{L}_{\sigma,s^{\prime}}$. We have
@@
Let $\mathcal{S}$ be a set of return morphisms from $\mathcal{A}^{*}$ to $\mathcal{A}^{*}$. The graph $\mathcal{G}^{L}(\mathcal{S})$ (resp., $\mathcal{G}^{R}(\mathcal{S})$) is defined by
@@
  • there is an edge from $T^{\prime}$ to $T$ labeled by $\sigma\in\mathcal{S}$ if $(T^{\prime},\sigma,T)$ is a left (resp., right) valid triplet.
@@
Let us prove that $G^{\prime}$ is a covering subgraph of $G^{L}(Y)$. The proof for $H^{\prime}$ and $G^{R}(Y)$ is similar. First, notice that $G^{L}(Y)$ and $G^{\prime}=\sigma^{L}(G)$ have the same set of vertices which is the image alphabet of $\sigma$. Moreover, they are acyclic for the coloring and multi-clique thus they are simple. Hence, it suffices to prove that if there is an edge in $G^{\prime}$, it is also an edge in $G^{L}(Y)$.
@@
Assume that $X$ is minimal dendric. The morphisms of $\mathcal{S}$ being return morphisms, the sequence ${\boldsymbol{\sigma}}$ is as in Theorem 42 and thus is primitive. Let us show that ${\boldsymbol{\sigma}}$ labels an infinite path in $\mathcal{G}^{L}(\mathcal{S})$. Similarly, it will label an infinite path in $\mathcal{G}^{R}(\mathcal{S})$. For all $n\geq 0$, as $X_{\boldsymbol{\sigma}}^{(n+1)}$ and $X_{\boldsymbol{\sigma}}^{(n)}$ are dendric and since $X_{\boldsymbol{\sigma}}^{(n)}$ is the image of $X_{\boldsymbol{\sigma}}^{(n+1)}$ under $\sigma_{n}$, the triplet $(G^{L}(X_{\boldsymbol{\sigma}}^{(n)}),\sigma_{n},G^{L}(X_{\boldsymbol{\sigma}}^{(n+1)}))$ is left valid by Proposition 38 and Proposition 41. Let $T_{0}$ be a covering tree of $G^{L}(X)=G^{L}(X_{\sigma}^{(0)})$. By Proposition 49, there exists a covering tree $T_{1}$ of $G^{L}(X_{\sigma}^{(1)})$ such that the triplet $(T_{0},\sigma_{0},T_{1})$ is left valid thus $\sigma_{0}$ labels an edge from $T_{0}$ to $T_{1}$ in $\mathcal{G}^{L}(\mathcal{S})$. We iterate the process to obtain an infinite path in $\mathcal{G}^{L}(\mathcal{S})$.

### v3

@@ 4. Other graphs characterizing dendricity
Let $X$ be the Thue-Morse shift space over $\{0,1\}$, i.e., the shift space generated by the substitution $\sigma:0\mapsto 01,1\mapsto 10$. The first few graphs $G^{L}_{n}(X)$ are given in Figure 2.
@@
Recall that a cycle or a path is simple if it uses distinct vertices (except for the first one and the last one in a cycle).
@@
Let $X$ be a shift space and $v\in\mathcal{L}_{n}(X)$.
The graph $\mathcal{E}_{X}(v)$ contains the path $(a_{1}^{L},b_{1}^{R},a_{2}^{L},\dots,b_{k}^{R},a_{k+1}^{L})$, with $a_{i}\neq a_{i+1}$, if and only if $G^{L}_{n+1}(X)$ contains the path
[…]
Symmetrically, the graph $\mathcal{E}_{X}(v)$ contains the path $(a_{1}^{R},b_{1}^{L},a_{2}^{R},\dots,b_{k}^{L},a_{k+1}^{R})$, with $a_{i}\neq a_{i+1}$, if and only if $G^{R}_{n+1}(X)$ contains the path
[…]
@@
Let $X$ be a shift, $N\geq 0$ and $C\subseteq\mathcal{A}$. If the graph $\mathcal{E}_{X}(v)$ is acyclic for all $v\in\mathcal{L}_{<N}(X)$, then the following properties are equivalent.
@@
Let us show the first set of equivalences. We denote by $H_{n}$ the subgraph of $G^{L}_{n}(X)$ generated by the vertices in $C$. Assume that the first property is satisfied. We prove that $H_{n}$ is connected by induction on $n$. The graph $G^{L}_{0}(X)$ is a complete graph thus any subgraph is connected. If $H_{n}$ is connected, then, to prove that $H_{n+1}$ is also connected, it suffices to show that any two vertices $a,b\in C$ that were connected by an edge in $H_{n}$ are connected by a path in $H_{n+1}$. Let $a,b\in C$ be two such vertices and $v$ be the label of an edge between them. As $a$ and $b$ are left extensions of $v$, $a^{L}$ and $b^{L}$ are connected in $\mathcal{E}_{X}(v)$ by a path avoiding vertices $c^{L}$, $c\notin C$. The vertices $a$ and $b$ are thus connected by a path in $H_{n+1}$ by Lemma 24. Note that the acyclic hypothesis is not needed for this implication.
@@ 5. Stabilization and eventually dendric shifts
To make the distinction clearer, we will talk about multi-graphs with colored edges instead of labeled edges. Formally, a multi-graph with colored edges is a multi-graph with labeled edges, but we use the terminology “colored” to highlight the fact that we are not interested in the label but in the sets of edges having the same label. We naturally adapt the terminology of acyclicity for the labeling to acyclicity for the coloring.
@@
A multi-graph $G$ with colored edges and with set of vertices $V$ is a multi-clique if there exist subsets $C_{1},\dots,C_{k}$ of $V$ such that the set of edges is the union of the sets of colored edges $\left((C_{i}\times C_{i})\setminus\operatorname{diag}(C_{i})\right)\times\{c_{i}\}$, $i\leq k$, where $c_{i}$ is the color of the edges and $c_{1},\dots,c_{k}$ are distinct colors. The multi-clique $G$ is then denoted $G(\{C_{1},\dots,C_{k}\})$.
@@
If $C=\{0,1,2,3\}$, $C_{1}=\{0,1\}$, $C_{2}=\{1,2,3\}$, $C_{3}=\{0,3\}$, the multi-clique $G(\{C_{1},C_{2},C_{3}\})$ is represented in Figure 4. It is not acyclic for the coloring because of the cycle $0,1,3,0$.
@@
Another way to see the graph $G^{L}(X)$ is to consider the right infinite words and their left extensions. More precisely, for a right infinite word $x$, we denote by $E^{L}_{X}(x)$ the set of letters $a$ such that $\mathcal{L}(ax)\subseteq\mathcal{L}(X)$. We then have
@@
Let $\sigma$ be the morphism defined by $\sigma(0)=0110$ and $\sigma(1)=011$. It is a return morphism for $01$. Let $X$ be the image under $\sigma$ of a Sturmian shift. It is not a dendric shift space since the graph $\mathcal{E}_{X}(\varepsilon)$ contains a cycle but, as the $\sigma$-initial factors are of length at most 4 and as every factor of a Sturmian shift space is ordinary, $X$ is eventually dendric with threshold at most 4 by Remark 15. We can then show that the only words which are not dendric are $\varepsilon$ and $1$ thus $G^{L}(X)=G^{L}_{2}(X)$ and $G^{R}(X)=G^{R}_{2}(X)$. These graphs are acyclic for the coloring and connected since $01$ (resp., $10$) is the only left (resp., right) special word of length 2.
@@ 6. Action of morphisms on $G^{L}(X)$ and $G^{R}(X)$
Let us prove the link between $G^{L}(X)$ and $G^{L}(Y)$. By Proposition 19, $Y$ is also eventually dendric. If we write $K=\max\{|\sigma(a)|:a\in\mathcal{A}\}$, let $N\in\mathbb{N}$ (resp., $M\in\mathbb{N}$) be large enough so that $N$ (resp., $M-K$) satisfies the condition 2 of Proposition 17 for the shift $X$ (resp., $Y$). Let us also assume that the antecedent of any word $u\in\mathcal{L}(Y)$ of length at least $M$ is of length at least $N$. By Proposition 32, $G^{L}(X)$ (resp., $G^{L}(Y)$) is constructed with the left special factors of length $N$ (resp., $M$).
We define the application $f$ mapping any left special word $u\in\mathcal{L}_{M}(Y)$ to the pair $(s_{u},v)\in\mathcal{T}^{L}(\sigma)\times\mathcal{L}_{N}(X)$ where $(s_{u},v_{u},p_{u})$ is the triplet associated with $u$ by Proposition 8 and $v$ is the prefix of length $N$ of $v_{u}$.
@@
$E^{L}_{Y}(u)\subseteq E^{L}_{Y}(s\sigma(v)w)=\varphi^{L}_{\sigma,s}(E^{L}_{X}(v))$
@@
The application $f$ is injective. Indeed, if $f(u)=f(u^{\prime})=(s,v)$, then $s_{u}=s=s_{u^{\prime}}$ and $v$ is a prefix of both $v_{u}$ and $v_{u^{\prime}}$. However, as $u$ and $u^{\prime}$ are left special, $v_{u}$ and $v_{u^{\prime}}$ must be left special. As they have a common prefix of length $N$ and $N$ is chosen large enough, it means that one is prefix of the other. Let us assume that $v_{u}$ is a prefix of $v_{u^{\prime}}$. Then $s\sigma(v_{u})w$ is a prefix of both $u$ and $u^{\prime}$. Since $u$ is of length $M$, $s\sigma(v_{u})w$ is a left special factor of length at least $M-K$ and, as $M$ is large enough, it is prefix of a unique left special factor of length $M$. This proves that $u=u^{\prime}$ and that $f$ is injective.
@@
$E^{L}_{Y}(u)\supseteq E^{L}_{Y}(s\sigma(v^{\prime})w)=\varphi_{\sigma,s}^{L}(E^{L}_{X}(v^{\prime}))=\varphi_{\sigma,s}^{L}(E^{L}_{X}(v)).$
@@ 7.1. $S$-adic representations
As $\max_{a\in\mathcal{A}_{N}}|\sigma_{[n,N)}(a)|$ goes to infinity when $N$ increases, $\mathcal{L}^{(n)}({{\boldsymbol{\sigma}}})$ defines a non-empty shift space $X_{{\boldsymbol{\sigma}}}^{(n)}$. More precisely, $X_{{\boldsymbol{\sigma}}}^{(n)}$ is the set of points $x\in\mathcal{A}_{n}^{\mathbb{Z}}$ such that $\mathcal{L}(x)\subseteq\mathcal{L}^{(n)}({{\boldsymbol{\sigma}}})$. Note that it may happen that $\mathcal{L}(X_{{\boldsymbol{\sigma}}}^{(n)})$ is strictly contained in $\mathcal{L}^{(n)}({{\boldsymbol{\sigma}}})$. Also observe that for all $n$, $X_{{\boldsymbol{\sigma}}}^{(n)}$ is the image of $X_{{\boldsymbol{\sigma}}}^{(n+1)}$ under $\sigma_{n}$.
We set $X_{{\boldsymbol{\sigma}}}=X_{{\boldsymbol{\sigma}}}^{(0)}$ and call ${\boldsymbol{\sigma}}$ an $S$-adic representation of $X_{{\boldsymbol{\sigma}}}$. If we want to specify that the morphisms $\sigma_{n}$, $n\geq 0$, belong to a given family $\mathfrak{S}$, we say that ${\boldsymbol{\sigma}}$ is an $\mathfrak{S}$-adic representation.
@@
We let $\mathcal{R}_{X}(w)$ denote the set of return words to $w$ in $X$ and we omit the subscript $X$ whenever it is clear from the context. The shift space $X$ being minimal, $\mathcal{R}(w)$ is finite (see [20]) thus we write $R_{X}(w)=\{1,\dots,\operatorname{Card}(\mathcal{R}_{X}(w))\}$. A morphism $\sigma:{R(w)}^{*}\to\mathcal{A}^{*}$ is a coding morphism associated with $w$ if $\sigma(R(w))=\mathcal{R}(w)$. It is trivially a return morphism for $w$.
Let us consider the set $\mathcal{D}_{w}(X)=\{x\in{R(w)}^{\mathbb{Z}}\mid\sigma(x)\in X\}$. It is a minimal shift space, called the derived shift of $X$ (with respect to $w$). We now show that derivation of minimal shift spaces allows us to build primitive $\mathfrak{S}$-adic representations with return morphisms. We inductively define the sequences $(w_{n})_{n\geq 0}$, $(R_{n})_{n\geq 0}$, $(X_{n})_{n\geq 0}$ and $(\sigma_{n})_{n\geq 0}$ by
@@
Conversely, if ${\boldsymbol{\sigma}}$ is an $S$-adic representation of a minimal shift space $X$ and ${\boldsymbol{\sigma}}$ contains only return morphisms, then one can choose words $w_{n}$ such that ${\boldsymbol{\sigma}}$ is defined as above. Hence, it is primitive.
@@
Let $X$ be a minimal dendric shift over the alphabet $\mathcal{A}=\{1,\dots,d\}$. For any $w\in\mathcal{L}(X)$, $\mathcal{D}_{w}(X)$ is a minimal dendric shift over $\mathcal{A}$ and the coding morphism associated with $w$ is a tame automorphism of $F_{\mathcal{A}}$. As a consequence, if ${\boldsymbol{\sigma}}=(\sigma_{n})_{n\geq 0}$ is the primitive $S$-adic representation from Theorem 42, then all morphisms $\sigma_{n}$ are tame dendric return morphisms.
@@
In particular, if $w_{n}$ is a letter, the morphism $\sigma_{n}$ is strongly left proper, i.e. the image of each letter begins with the letter $w_{n}$ which does not appear elsewhere. Moreover, if $w_{n}$ is a bispecial letter, then $w_{n}\in\mathcal{T}^{R}(\sigma_{n})$ and $\varepsilon\in\mathcal{T}^{L}(\sigma_{n})$. We can always chose such a $w_{n}$ as shown in the following results. The first one is a trivial lemma relying on the minimality of $X$.
@@
First, note that, because $\mathcal{E}_{X}(\varepsilon)$ is connected, there is always a left and a right special letter. Let $a_{1}$ be a right special letter. If it is bispecial, we can conclude. Otherwise, let $a_{2}$ be its unique left extension. Using Lemma 44, we know that $a_{2}\neq a_{1}$. In the graph $\mathcal{E}_{X}(\varepsilon)$, $a_{2}^{L}$ is then the only neighbour of the vertex $a_{1}^{R}$ but this graph is connected thus the vertex $a_{2}^{L}$ is of degree at least two and $a_{2}$ is a right special letter. Once again, if it is left special, we have found a bispecial letter and otherwise it has a unique left extension $a_{3}$. By Lemma 44, we must have $a_{3}\neq a_{1}$ and $a_{3}\neq a_{2}$. We can iterate the process to define $a_{4},a_{5},\dots$. However, since the alphabet is finite, it stops at some point, meaning that we have found a bispecial letter. ∎
@@ 7.2. Characterization via two graphs
We can deduce that a shift space is minimal dendric if and only if it has a primitive $S$-adic representation labeling an infinite path in the graphs built with the left (resp., right) valid triplets for dendric return morphisms. The details are not given here as the idea is the same as in the stronger version that we will prove. Indeed, instead of considering all the valid triplets, we can restrict ourselves to the case where the graphs are colored trees, i.e. all the edges have a different color and the corresponding uncolored graph is a tree.
Note that a colored tree is trivially acyclic for the coloring and connected. As we only work with multi-cliques, in what follows, any colored graph whose underlying graph is a tree is a colored tree. Conversely, an uncolored tree corresponds to a unique colored tree so we will make no distinction between the colored and the uncolored version.
@@
Recall that a covering tree of a graph $G$ is a subgraph $T$ of $G$ with the same vertices as $G$ and which is a tree. If $G$ is a colored graph, then $T$ does not need to share the same edges colors as $G$ and is always assumed to be a colored tree (or an uncolored tree, as explained above).
@@
  • Otherwise, $as$ and $s^{\prime}$ are not suffix comparable thus $\mathcal{A}^{L}_{\sigma,s^{\prime}}\cap C_{i}=\mathcal{A}^{L}_{\sigma,s^{\prime}}\cap E$, $\mathcal{A}^{L}_{\sigma,s^{\prime}}\cap D\subseteq\{c\}$ and no edge was lost from $G^{L}_{\sigma,s^{\prime}}$ to $H^{L}_{\sigma,s^{\prime}}$. We have
@@
Let $\mathfrak{S}$ be a set of return morphisms from $\mathcal{A}^{*}$ to $\mathcal{A}^{*}$. The graph $\mathcal{G}^{L}(\mathfrak{S})$ (resp., $\mathcal{G}^{R}(\mathfrak{S})$) is defined by
@@
  • there is an edge from $T^{\prime}$ to $T$ labeled by $\sigma\in\mathfrak{S}$ if $(T^{\prime},\sigma,T)$ is a left (resp., right) valid triplet.
@@
Let us prove that $G^{\prime}$ is a covering subgraph of $G^{L}(Y)$. The proof for $H^{\prime}$ and $G^{R}(Y)$ is similar. First, note that $G^{L}(Y)$ and $G^{\prime}=\sigma^{L}(G)$ have the same set of vertices which is the image alphabet of $\sigma$. Moreover, they are acyclic for the coloring and multi-clique thus they are simple. Hence, it suffices to prove that if there is an edge in $G^{\prime}$, it is also an edge in $G^{L}(Y)$.
@@
Assume that $X$ is minimal dendric. The morphisms of $\mathfrak{S}$ being return morphisms, the sequence ${\boldsymbol{\sigma}}$ is as in Theorem 42 and thus is primitive. Let us show that ${\boldsymbol{\sigma}}$ labels an infinite path in $\mathcal{G}^{L}(\mathfrak{S})$. Similarly, it will label an infinite path in $\mathcal{G}^{R}(\mathfrak{S})$. For all $n\geq 0$, as $X_{\boldsymbol{\sigma}}^{(n+1)}$ and $X_{\boldsymbol{\sigma}}^{(n)}$ are dendric and since $X_{\boldsymbol{\sigma}}^{(n)}$ is the image of $X_{\boldsymbol{\sigma}}^{(n+1)}$ under $\sigma_{n}$, the triplet $(G^{L}(X_{\boldsymbol{\sigma}}^{(n)}),\sigma_{n},G^{L}(X_{\boldsymbol{\sigma}}^{(n+1)}))$ is left valid by Proposition 38 and Proposition 41. Let $T_{0}$ be a covering tree of $G^{L}(X)=G^{L}(X_{\sigma}^{(0)})$. By Proposition 49, there exists a covering tree $T_{1}$ of $G^{L}(X_{\sigma}^{(1)})$ such that the triplet $(T_{0},\sigma_{0},T_{1})$ is left valid thus $\sigma_{0}$ labels an edge from $T_{0}$ to $T_{1}$ in $\mathcal{G}^{L}(\mathfrak{S})$. We iterate the process to obtain an infinite path in $\mathcal{G}^{L}(\mathfrak{S})$.